OpenStudy (anonymous):

Explain, in complete sentences, how you would use the two methods (distribution and box method) to find the product of (x – 2) (x – 3). Please include your steps for each method and the resulting product. Which method do you prefer and why? I need help with the box method part. Idk how to do it on the computer?? Can someone help me with this?? Please and thank you :)

OpenStudy (anonymous):

This is what I have so far: I prefer the distribution method because, first of all, I find it easier, and secondly, that is how Mrs. Huerkamp explained it in the live lesson. To do this you simply utilize the FOIL technique by multiplying the first, outer, inner and last variables together. (x) (x) = x2 (x) (-3) = -3x (-2) (x) = -2x (-2) (-3) = 6 so your answer is: x2 - 3x - 2x + 6 = x2 - 5x + 6 The box method would be you make a 2x2 box and go up, down, and sideways to multiply the variables. It’s the same basic principal, just a little different: